Write the equation for the following relation.
Q = {(x, y): (2, 8), (3, 27), (4, 64), (5, 125), . . .}

Answer by Edwin McCravy(20086)

  (2,8) = (2,2·2·2) = (2,2³)
 (3,27) = (3,3·3·3) = (3,3³)
 (4,64) = (4,4·4·4) = (4,4³)
(5,125) = (5,5·5·5) = (5,5³)
  (x,y) = (x,x·x·x) = (x,x³)

Therefore y = x³

Edwin